OpenStudy (baby456):

5q-7=2/3 help me please

OpenStudy (anonymous):

What are we doing here? Simplifying or?

OpenStudy (baby456):

simplyfiying

OpenStudy (anonymous):

We simplify the equation to the form, which is simple to understand 5q-7=2/3

OpenStudy (anonymous):

Correct?

OpenStudy (anonymous):

Simplifying: 5q-7=0.67

OpenStudy (anonymous):

We move all terms containing q to the left and all other terms to the right. +5q=0.67+7

OpenStudy (baby456):

ok

OpenStudy (anonymous):

We simplify left and right side of the equation. +5q=+7.67

OpenStudy (anonymous):

Do you know what to do next?

OpenStudy (baby456):

q=?

OpenStudy (baby456):

12?

OpenStudy (anonymous):

Close. Not quite, take a look at the steps again.

OpenStudy (anonymous):

What do you do next to get the solution?

OpenStudy (baby456):

1.

OpenStudy (anonymous):

We divide both sides of the equation by 5 to get q. q=1.53

OpenStudy (anonymous):

Do you understand?

OpenStudy (baby456):

how

OpenStudy (baby456):

no

OpenStudy (anonymous):

Oh wait, the answer is 12. Im sorry, i missed calculated:)

OpenStudy (baby456):

ok it okAY

OpenStudy (baby456):

THANKS YOU

OpenStudy (anonymous):

No problem:) happy to help you:)